OpenStudy (anonymous):

Solve the following system of equations using elimination: 4x - 2y = 8 3x - 9y = -54

OpenStudy (danjs):

multiply the first equation by -3 and the second equation by 4 then add them

OpenStudy (danjs):

-3(4x-2y) = -3*8 4( 3x-9y) = 4* -54

OpenStudy (danjs):

the x term will cancel when you add them together, then solve for y, Use that y value in any of the 2 equations to find x